OpenStudy (anonymous):

Darlene Fine wants to have at least $50,000 in her savings account in 10 years. If her account pays 3.6% interest compounded annually, what should Darlene's initial investment be if she plans to keep the account without making deposits or withdrawals? Hint: Use the formula for annual compounding. $36,764.71 $14,894.72 $16,764.71 $35,105.28

OpenStudy (anonymous):

A=P(1+r100)n A-amount of money accumulated after n years, including interest. P-principal amount (the initial amount you borrow or deposit) r-annual rate of interest n-number of years So you know that A=$50,000 r=3.6 and n=10,so: 50,000=P(1+0.036)10

OpenStudy (anonymous):

hello baby

undeadknight26 (undeadknight26):

^ ignore the thirsty guy.

OpenStudy (anonymous):

lol

undeadknight26 (undeadknight26):

Are you familiar with: 50,000=P(1.36)^10?

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

sorta

undeadknight26 (undeadknight26):

Pretty much you solve that and you will get your answer :D

undeadknight26 (undeadknight26):

My bad type: 50,000=P(1.036)^10

OpenStudy (anonymous):

mhmm

undeadknight26 (undeadknight26):

lol so what do you get O.o

OpenStudy (anonymous):

well idk what P is

OpenStudy (anonymous):

2.84

undeadknight26 (undeadknight26):

Well your solving for P...So you would do: 50000=P(1.036)^10 50000 = P(1.036)^10 50000 = p(1.424287143492311) 50000/1.424287143492311 = P

undeadknight26 (undeadknight26):

So after all that hard shtupid work ._. Your answer is ???

OpenStudy (anonymous):

35105.2807?

undeadknight26 (undeadknight26):

YOSHHHHHHHHHH <3 So whats the answer to your question ma'am?

OpenStudy (anonymous):

thyank you :) I understand now mm yoiu think you can hellp me with this one ? Joshua Treet had a $400.00 average monthly expenditure for entertainment during the second quarter of the year. He spent $401.50 in July and $250.00 in August. How much can Joshua spend in September to have the same average as in the second quarter?

OpenStudy (anonymous):

2Q: $400 monthly for Entertainment July: $401.50 August:$250

OpenStudy (anonymous):

September = ?

OpenStudy (anonymous):

idk how to work this if i add or multiply or what

undeadknight26 (undeadknight26):

Google is my best friend....It always hints me off at stuff!

OpenStudy (anonymous):

can u help with this one?

OpenStudy (anonymous):

i got 548

undeadknight26 (undeadknight26):

Show ya work :D

OpenStudy (anonymous):

well i added july and august $ and 548 then divided by 3 got 400

undeadknight26 (undeadknight26):

Sounds right =.= Is it right? Maybe...Or maybe Not o.O

OpenStudy (anonymous):

its right

undeadknight26 (undeadknight26):

Good job :D